OpenStudy (anonymous):

A recursive function is shown below: f(1) = 5 and f(n) = f(n - 1) - 6; n > 1 Which of the following lists the terms in the sequence defined by this recursive function?

OpenStudy (anonymous):

1) 5, 1, 7, 13, 19, ... 2) 5, -1, -7, -13, -19, ... 3) 5, 11, 17, 23, 29, ... 4) 5, -11, -17, -23, -29, ...

OpenStudy (anonymous):

If you're given \(f(1)=5\), then by the definition of the recursive function, we see that \[\begin{aligned} f(2) &= f(1) - 6 = \ldots\\ f(3) &= f(2)-6 = \ldots\\ f(4) &= f(3) - 6 =\ldots\end{aligned}\] Can you proceed from here to see what values you get?

OpenStudy (anonymous):

Yes, thank you